Please show all steps!!! The answer is 69.8 torr.

16. A solution is made by adding 0.100 mole of ethyl ether to 0.628 mole of ethyl alcohol. If the vapor pressure of ethyl ether and ethyl alcohol at 20°C are 375 torr and 20.0 torr, respectively.  What is the vapor pressure of the solution at 20°C (assuming ideal behavior)?
Answer #1

Mole fraction = mole of A/(mole of A+mole of B)

Mole fraction of ethyl ether= 0.100/(0.100+0.628)=0.1374

Mole fraction of ethyl alcohol= 0.628/(0.100+0.628)=0.863

Vapour pressure=(0.1374*375)+(0.863*20)= 68.8 Answer